Given the information


Product.........Qty produced ......price...........Qty produced ........ Price .............................(kg) 1995............... 1995. ................(kg)......,..............2000


Maize...................50 .,........................N20................. 100 ................. N30

Rice........................80 ........................ N40 ................. 160 ................... N60



Calculate:


The GNP at market 2000 prices, GNP at current 1995 prices GNP at constant 1995 prices in year 2000

The GNP price deflator (Nominal GNP)/Real GNP


1
Expert's answer
2021-07-12T01:06:03-0400

i.

GNP = Quantity "\times" price year

"1995 \\ (50\times 20) + (80\times 40) = N4200\\

2000 \\ (1000\times 30) + (160\times 60) =N12600\\"


ii.

GNP for 2000 year using 1995 as the base year.

GNP = Quantity "\times" price of the base year

"(1000\times 20) + (160\times 40) =N8400\\"


iii.

GNP deflator = [GNP for 2000 "\div" GNP for 2000 year using 1995 as the base year]"\times 100"

"2000\\ (\frac{12600 } {8400}) \times 100 = 150\\"